Question 25
ProbabilityThe probability that Olu, Ade and shola will hit a target are 2/3, 3/4 and 4/5 respectively. Find the probability that only Olu will hit the target
- A.7/60
- B.1/30
- C.1/60
- D.2/5
Show the answer and explanation
Answer: B (1/30)
Explanation
Olu hits and the other two miss: 2/3 × 1/4 × 1/5 = 2/60 = 1/30.